# VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents

### v2

@@ A.5 Complete Results Of Text-Speech Alignment
GLM4-Voice 65.00 (-17.20) 82.98 (-1.28) 47.50 (-32.50) 82.40 (0.20) 85.00 (-0.62) 82.67 (-2.66) 83.20 (3.20) 79.44 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
Kimi-Audio 45.60 (-30.80) 65.96 (-7.23) 25.83 (-53.74) 56.20 (-20.60) 54.38 (-23.74) 77.33 (18.66) 73.60 (-7.20) 36.67 (-43.89) 93.80 (-1.60) 93.20 (-1.90) 94.40 (-1.20) 87.80 (-4.40)
Qwen2.5-Omni-7B 75.60 (-5.40) 82.55 (1.27) 68.75 (-12.08) 77.80 (-3.40) 79.38 (-1.87) 81.33 (5.33) 74.40 (-8.00) 77.22 (-5.56) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
Baichuan-Audio-Chat 60.40 (-0.80) 61.70 (-0.43) 59.17 (-1.26) 61.60 (-0.60) 61.88 (-0.62) 62.67 (-5.33) 68.00 (3.20) 56.67 (-1.11) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
Qwen2-Audio-7B-Instruct 76.40 (-5.80) 80.43 (-1.35) 72.50 (-10.00) 81.60 (0.80) 79.38 (0.63) 80.00 (-6.67) 83.20 (0.00) 83.33 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
StepAudio 62.53 (-18.94) 77.87 (-5.53) 47.50 (-32.08) 71.48 (-3.71) 68.75 (3.75) 76.00 (-13.33) 74.40 (-0.80) 70.00 (-8.33) 95.80 (1.00) 96.00 (1.20) 95.60 (0.80) 38.26 (-54.78)
StepAudio2 75.60 (-4.80) 75.32 (-2.13) 75.83 (-7.50) 75.40 (-2.00) 78.75 (-0.63) 88.00 (2.67) 67.20 (-4.80) 72.78 (-3.33) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
Mimo-Audio 76.00 (-14.80) 92.77 (-0.42) 59.58 (-28.75) 88.40 (-0.20) 90.62 (2.50) 96.00 (10.67) 88.80 (1.60) 82.78 (-8.33) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
GPT4o-Audio 79.00 (-8.60) 88.94 (0.00) 69.17 (-17.08) 88.40 (2.20) 91.88 (4.38) 89.33 (0.00) 85.60 (0.80) 86.67 (2.23) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
Qwen3-Omni 85.47 (-9.48) 96.60 (1.71) 74.58 (-20.42) 95.19 (1.12) 93.12 (3.74) 98.67 (2.67) 96.80 (0.80) 94.44 (-1.67) 98.80 (-0.80) 98.80 (-0.80) 98.80 (-0.80) 95.22 (-3.91)
@@
GLM4-Voice 64.00 (-22.60) 45.00 (-36.75) 96.57 (-0.57) 74.86 (-12.57) 89.20 (-0.80) 85.56 (-3.89) 82.94 (-1.47) 93.03 (-0.68) 92.82 (1.57) 78.40 (-14.60)
Kimi-Audio 74.00 (-5.40) 60.50 (-11.65) 94.86 (4.57) 84.00 (-1.14) 75.80 (-9.60) 76.73 (-8.36) 58.24 (-11.47) 93.71 (0.45) 70.25 (-19.75) 76.80 (-11.20)
Qwen2.5-Omni-7B 69.40 (-16.80) 52.00 (-30.50) 95.43 (0.57) 82.86 (-2.85) 85.20 (-4.80) 80.74 (-6.17) 74.71 (-8.23) 95.51 (-1.79) 85.57 (-4.43) 79.80 (-10.20)
Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.14 (-8.57) 58.29 (-12.57) 73.40 (-2.80) 73.45 (0.36) 64.41 (0.88) 81.35 (-5.39) 72.00 (-5.50) 67.20 (-4.60)
Qwen2-Audio-7B-Instruct 59.80 (-17.60) 37.75 (-32.00) 93.71 (-1.72) 76.00 (-1.14) 87.60 (-1.60) 82.18 (-6.18) 80.00 (-1.49) 96.18 (-0.18) 88.50 (0.25) 77.80 (-13.40)
StepAudio 55.87 (-30.80) 36.00 (-46.25) 93.71 (-0.58) 63.43 (-25.71) 86.92 (-2.94) 83.27 (-10.18) 77.65 (-1.47) 97.75 (0.45) 85.25 (-3.00) 66.50 (-27.25)
StepAudio2 65.00 (-19.20) 45.75 (-33.00) 94.86 (-1.14) 78.86 (-6.28) 83.40 (-5.80) 84.00 (-2.18) 66.18 (-12.06) 94.83 (-2.47) 84.75 (-6.75) 81.60 (-8.40)
Mimo-Audio 73.20 (-20.20) 54.00 (-35.50) 99.43 (4.43) 90.86 (-4.57) 90.40 (-5.00) 86.91 (-9.09) 86.76 (-5.00) 99.55 (0.45) 85.50 (-8.25) 95.20 (-4.20)
GPT4o-Audio 68.20 (-19.20) 61.00 (-20.50) 98.29 (-0.57) 54.86 (-34.28) 92.00 (-1.60) 86.18 (-6.55) 87.94 (-0.30) 96.40 (0.22) 94.25 (-1.50) 74.60 (-21.60)
Qwen3-Omni 83.47 (-13.86) 71.25 (-25.50) 97.71 (-1.72) 97.14 (0.57) 96.44 (-1.85) 96.73 (-1.45) 94.71 (-2.94) 100.00 (0.00) 93.75 (-3.25) 96.25 (-3.75)
@@
GLM4-Voice 84.41 (-6.90) 85.81 (-4.04) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
Kimi-Audio 77.71 (-9.40) 81.18 (-4.44) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
Qwen2.5-Omni-7B 80.44 (-6.75) 83.25 (-2.17) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
Baichuan-Audio-Chat 67.46 (-7.02) 68.57 (-5.52) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
Qwen2-Audio-7B-Instruct 79.60 (-9.04) 87.29 (-2.56) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
StepAudio 85.81 (-5.64) 86.21 (-3.35) 89.44 (-2.69) 90.43 (-2.32) 82.42 (-7.78) 81.30 (-13.90)
StepAudio2 83.04 (-4.79) 85.02 (-2.27) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
Mimo-Audio 95.10 (-3.04) 96.45 (-0.69) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
GPT4o-Audio 92.34 (-1.94) 92.12 (-0.29) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)
Qwen3-Omni 97.61 (-1.62) 98.92 (-0.19) 94.61 (-4.04) 95.36 (-3.77) 96.92 (-2.42) 100.00 (0.00)
@@
[…]
GLM4-Voice 79.2 (83.2) 80.8 (88.4) 78.8 (81.8) 88.6 (92.4) 84.0 (88.4) 75.4 (75.6) 88.2 (92.6)
Kimi-Audio 75.4 (76.8) 81.4 (86.6) 64.2 (64.8) 81.0 (84.8) 83.2 (88.4) 75.8 (76.6) 82.0 (83.4)
Qwen2.5-Onmi 71.4 (72.4) 75.4 (80.4) 74.2 (75.8) 84.8 (86.2) 72.8 (79.0) 80.0 (80.8) 77.8 (79.2)
Baichuan-Audio-Chat 83.2 (81.8) 86.0 (89.2) 73.6 (71.4) 85.6 (90.2) 75.2 (79.4) 78.6 (78.8) 92.6 (93.4)
StepAudio2Mini 77.0 (82.2) 77.4 (86.6) 70.8 (73.8) 79.4 (88.0) 81.8 (87.8) 62.6 (63.8) 83.8 (95.2)
GPT4o-Audio 86.6 (89.2) 85.4 (93.4) 83.2 (85.6) 91.2 (96.6) 87.4 (93.4) 83.6 (85.8) 95.2 (98.2)
Table 20: A2A Result in TIF - A2A W/ ASR (A2A W/O ASR)
[…]
GLM4-Voice 65.0 (81.6) 75.4 (83.8) 72.6 (77.4) 78.0 (85.4) 77.8 (83.8) 69.2 (76.6) 86.0 (89.8)
Kimi-Audio 50.4 (62.0) 60.4 (83.4) 53.0 (56.0) 62.4 (80.8) 65.6 (79.0) 78.4 (83.4) 58.4 (69.8)
Qwen2.5-Omni 57.0 (59.6) 58.2 (64.2) 56.4 (60.2) 66.4 (72.4) 56.2 (66.6) 73.4 (75.6) 56.2 (63.0)
Baichuan-Audio-Chat 79.8 (77.6) 80.0 (81.0) 75.8 (75.0) 77.0 (76.8) 79.8 (81.2) 72.0 (72.0) 86.2 (86.0)
StepAudio2Mini 66.2 (67.0) 73.0 (78.0) 55.2 (56.2) 75.0 (77.0) 70.6 (73.0) 69.6 (70.4) 84.2 (87.8)
GPT4o-Audio 90.2 (91.4) 90.0 (94.8) 85.0 (85.8) 93.4 (95.8) 92.4 (94.6) 80.4 (81.6) 96.0 (98.4)
Table 21: A2A Result in TIF-En - A2A W/ ASR (A2A W/O ASR)

### v3

@@ A.4 Complete Results Of Robustness
GLM-4-Voice 63.58 (-16.21) 81.28 (2.13) 46.25 (-34.17) 79.26 (0.56) 81.25 (0.63) 78.67 (-5.33) 81.60 (3.20) 76.11 (1.11) 93.00 (-1.20) 91.60 (-2.00) 94.40 (-0.40) 39.13 (-49.13)
Kimi-Audio 45.89 (-24.43) 62.98 (-5.53) 29.17 (-42.91) 45.74 (-20.00) 35.62 (-30.00) 45.33 (-1.34) 65.60 (-4.00) 41.11 (-30.00) 92.00 (-3.00) 92.80 (-3.60) 91.20 (-2.40) 86.09 (-3.91)
Qwen2.5-Omni 74.53 (1.06) 75.32 (2.55) 73.75 (-0.42) 75.74 (-0.56) 77.50 (-0.62) 80.00 (0.00) 76.80 (4.80) 71.67 (-4.44) 82.60 (-2.20) 82.80 (-1.60) 82.40 (-2.80) 79.13 (0.00)
Baichuan-Audio 71.58 (-3.16) 74.89 (-3.83) 68.33 (-2.50) 77.96 (4.81) 75.62 (0.00) 84.00 (8.00) 80.00 (7.20) 76.11 (6.11) 90.40 (-2.20) 89.20 (-3.20) 91.60 (-1.20) 64.78 (-16.96)
Step-Audio 2 mini 69.26 (-5.69) 69.36 (-5.96) 69.17 (-5.41) 69.63 (-1.85) 75.62 (-1.26) 77.33 (12.00) 61.60 (-6.40) 66.67 (-5.00) 89.60 (0.40) 90.40 (1.20) 88.80 (-0.40) 69.13 (-13.91)
MiMo-Audio 85.26 (-1.90) 87.23 (0.00) 83.33 (-3.75) 84.07 (-1.12) 78.75 (-6.87) 89.33 (2.66) 85.60 (0.00) 85.56 (1.67) 97.20 (-2.00) 96.00 (-3.60) 98.40 (-0.40) 50.43 (-41.31)
GPT-4o-Audio 77.89 (-5.69) 87.23 (4.25) 68.75 (-15.42) 85.74 (1.85) 85.00 (-1.25) 86.67 (6.67) 88.00 (2.40) 84.44 (2.22) 92.80 (-2.00) 92.80 (-1.20) 92.80 (-2.80) 45.65 (-46.96)
Qwen3-Omni 77.05 (-12.00) 89.36 (-0.85) 65.00 (-22.92) 87.59 (0.18) 84.38 (3.76) 92.00 (4.00) 90.40 (-3.20) 86.67 (-2.22) 95.80 (0.20) 94.80 (-1.60) 96.80 (2.00) 93.91 (-3.05)
Fun-Audio-Chat 86.11 (-0.63) 89.36 (2.13) 82.92 (-3.33) 86.85 (-0.37) 88.12 (1.24) 90.67 (0.00) 86.40 (-2.40) 84.44 (-0.56) 94.00 (-0.41) 94.00 (-0.47) 94.00 (-0.35) 86.09 (-9.13)
@@
GLM-4-Voice 61.73 (-20.14) 43.75 (-35.25) 94.86 (1.72) 69.71 (-7.43) 85.96 (-1.99) 80.36 (-5.82) 78.53 (-4.12) 91.69 (0.90) 89.75 (-0.75) 76.25 (-7.75)
Kimi-Audio 64.27 (-11.73) 47.75 (-22.00) 92.00 (3.43) 74.29 (-3.42) 66.51 (-12.46) 71.27 (-8.73) 37.35 (-24.41) 85.17 (-0.67) 67.25 (-18.00) 61.75 (-20.75)
Qwen2.5-Omni 60.80 (-8.93) 47.75 (-18.25) 81.14 (0.57) 70.29 (2.86) 80.14 (0.48) 81.09 (-1.46) 73.24 (-0.29) 82.92 (1.57) 82.25 (1.25) 79.00 (-7.75)
Baichuan-Audio 64.93 (-14.00) 48.50 (-24.75) 94.29 (2.29) 73.14 (-5.72) 85.48 (-0.89) 78.91 (-5.09) 78.24 (1.77) 94.61 (-1.12) 86.00 (0.00) 72.75 (-12.75)
Step-Audio 2 mini 55.47 (-22.66) 38.00 (-38.00) 81.14 (-4.57) 69.71 (-5.72) 77.33 (-3.56) 76.00 (-2.91) 66.76 (-8.24) 85.62 (0.23) 78.00 (-4.25) 74.25 (-3.25)
MiMo-Audio 73.07 (-15.60) 57.00 (-28.25) 99.43 (-0.57) 83.43 (-1.71) 91.78 (-1.78) 89.45 (-3.64) 85.29 (-5.00) 97.53 (0.00) 92.50 (0.25) 85.00 (-11.75)
GPT-4o-Audio 67.07 (-18.80) 60.75 (-24.00) 94.86 (2.29) 53.71 (-28.00) 88.56 (-1.51) 79.64 (-8.72) 86.76 (1.17) 93.03 (2.02) 91.25 (-2.75) 72.50 (-17.75)
Qwen3-Omni 78.53 (-14.67) 67.25 (-27.00) 90.86 (-4.00) 92.00 (2.86) 89.66 (-3.29) 84.73 (-2.54) 82.06 (-7.65) 99.10 (0.00) 89.00 (-3.75) 84.25 (-7.00)
Fun-Audio-Chat 70.13 (-20.94) 53.25 (-37.25) 96.00 (0.00) 82.86 (-4.57) 91.03 (-1.50) 82.91 (-4.36) 87.35 (-2.36) 96.40 (-1.35) 93.75 (1.50) 83.25 (-9.00)
@@
GLM-4-Voice 78.60 (-7.91) 80.79 (-5.42) 78.43 (-9.21) 82.90 (-4.06) 78.02 (-6.38) 71.30 (-16.53)
Kimi-Audio 71.51 (-10.04) 70.15 (-8.77) 65.84 (-16.86) 71.01 (-10.73) 74.07 (-7.91) 77.83 (-7.82)
Qwen2.5-Omni 78.27 (-2.83) 81.38 (0.20) 74.61 (-6.06) 78.55 (-0.58) 74.07 (-7.47) 78.91 (-3.48)
Baichuan-Audio 78.71 (-4.64) 83.74 (0.19) 73.48 (-11.01) 80.00 (-1.45) 72.09 (-7.25) 78.26 (-8.91)
Step-Audio 2 mini 74.78 (-4.78) 76.45 (-1.48) 73.48 (-6.30) 75.36 (-5.22) 70.11 (-7.47) 76.52 (-7.61)
MiMo-Audio 89.38 (-4.30) 92.12 (-1.48) 89.89 (-2.92) 91.01 (-2.61) 84.18 (-9.01) 86.74 (-8.48)
GPT-4o-Audio 86.51 (-1.98) 87.88 (-0.30) 85.84 (-3.37) 80.00 (-6.09) 84.18 (-1.97) 91.30 (-1.31)
Qwen3-Omni 86.03 (-2.32) 91.13 (0.39) 82.70 (-3.37) 83.77 (-6.09) 83.30 (-4.61) 82.39 (-2.18)
Fun-Audio-Chat 85.15 (-4.63) 88.87 (-1.08) 83.15 (-4.94) 84.06 (-4.35) 82.42 (-6.59) 82.39 (-10.44)